### v1

@@ 5.1 Hardware and Software
The experiments were conducted on two clusters. The first one, MarenostrumV [^2] https://www.bsc.es/supportkc/docs/MareNostrum5/overview (MNV), used 32 nodes from the general queue. Each node has two 56-core Intel Xeon 8480 CPUs (3584 cores in total). This system was employed to analyze the behaviour of both spawning techniques described in Section 4.
@@
The second cluster, Nasp [^3] https://www.hpca.uji.es/computing-facilities/, consists of two sets of 8 nodes. The first set has nodes with two 10-core Intel Xeon 4210 CPUs (160 cores total), connected through a 100 Gbps InfiniBand EDR and a 10 Gbps Ethernet network. The second set uses nodes with 32-core Intel Xeon 6346 CPUs (256 cores total) connected via 10 Gbps Ethernet. Both sets communicate through a shared 10 Gbps Ethernet link between their switches. This heterogenous cluster was used to evaluate the effectiveness of the improved method in Section 4.2.
@@
All experiments in MNV were run with MPICH 4.2.0 and CH4:OFI (InfiniBand), while in NASP with MPICH 3.4.3 and CH3:Nemesis (Ethernet). Both the Proteo version [^4] https://lorca.act.uji.es/gitlab/martini/malleability_benchmark/-/tree/Paper-ParallelSpawn and the experimental results [17] are publicly available.
Proteo was used to execute the different approaches. Before reconfiguration, 5 iterations of Pi computation with MPI_Allgather were performed to ensure MPI initialization. MaM was configured with Baseline (B) and Merge (M), each with both strategies described in 4, or Merge without strategies. For shrinking, only Baseline plus the described techniques was compared against a stand-alone Merge. Each configuration was executed 20 times, and the median result was reported for every triplet of sources, targets, and method.
5.2 Homogenous System Evaluation
@@
On the homogeneous cluster MNV, both strategies described in Section 4 were evaluated. In total, 5 configurations were tested for expansion and 3 for shrinking. Each configuration involved a single reconfiguration from $NS$ to $NT$ processes, using up to 42 combinations of nodes from the set 1,2,4,8,16,24,32{1,2,4,8,16,24,32}.
@@
Figure 4 reports the median reconfiguration times depending on the number of sources ($NS$) and targets ($NT$). The top Figure 4(a) shows expansion results, where the Merge method outperforms all others in 17 out of 21 cases ($80.9\%$). Parallel Merge follows closely, requiring up to $1.13\times$ more time due to additional synchronizations among groups. This overhead grows when more than 8 groups are created and the number of groups is not a power of two. In those cases, at least three steps are required for the binary connection, and the presence of unbalanced leaves increases the cost.
@@
Parallel Baseline methods are consistently slower (up to $1.73\times$), except in one case (1 node to 16 nodes). This overhead stems from spawning extra processes, which in addition leads to oversubscription.
@@
The bottom Figure 4(b) shows shrinking results. Shrinking without spawning processes is clearly advantageous, completing with a speedup of $1387\times$ over the other evaluated approaches. These findings are consistent with previous work [20].
@@
Figure 4: Resize times in homogenous system.
Figure 5 summarizes the best method for each ($NS$, NTNT) pair based on statistical tests. Axes are defined with NS on the vertical and NT on the horizontal; thus, the upper triangle corresponds to expansion and the lower triangle to shrinking. When multiple methods appear in a cell, they are statistically equivalent, ordered by ascending time.
The main difference arises in expansions with 8 or fewer groups, requiring at most 3 steps. In such cases, the parallel methods are preferred as they enable shrinking without spawning processes, while maintaining times comparable to Merge, though statistically higher.
Figure 5: Median reconfiguration times in MNV.
5.3 Improved Method Evaluation
On the heterogeneous cluster NASP, only the improved strategy from Section 4.2 was evaluated. Three configurations were tested for expansion and two for shrinking. Each configuration involved a single reconfiguration from $NS$ to $NT$ processes, using up to 72 combinations of nodes from the set 1,2,4,6,8,10,12,14,16{1,2,4,6,8,10,12,14,16}. In all cases, the number of nodes of each type was balanced (half of one type and half of the other). When only one node was used, the 20-core node was selected.
Figure 6 shows the median reconfiguration times as a function of $NS$ and $NT$. The top Figure 6(a) reports expansion results, where the Merge method consistently outperforms the others. The iterative diffusive variant of Merge achieves comparable results, with at most a $1.25\times$ increase, but its cost grows with the number of generated groups.
Baseline methods are always more expensive, ranging from $1.76\times$ to $4.92\times$ slower, due to oversubscription caused by spawning more processes.
The bottom Figure 6(b) shows shrinking results. As in the previous evaluation, Merge is the preferred method with a minimum speedup of $20\times$. To ensure it can be applied, a parallel spawn is required. This approach also benefits future shrink operations, adding only a minor overhead during expansion. However, in difference to the previous evaluation, in 28 out of 32 cases ($87.5\%$), standalone Merge delivered statistically lower times than its iterative diffusive variant.
@@ 5.3 Improved Method Evaluation
Figure 6: Resize times in heterogenous system.
@@ 6 Conclusions and Future Work
This work has introduced and evaluated two new strategies for spawning processes in the traditional MPI. Both approaches rely on parallel spawning with two main goals: (1) reducing the cost of process creation, and (2) enabling efficient shrink operations with the Merge method, the fastest available option. Achieving the second goal requires for each MCW to be fully contained within a node. Previously, this was not possible: when shrinking, processes that should be terminated remained as zombies because other ranks of the same MCW were still active on different nodes, preventing node reuse. The proposed strategies resolve this issue, enabling fast shrink operations.
Between the two strategies, the Hypercube approach is limited to homogeneous, non-shared systems, while the Iterative Diffusive strategy also applies to heterogeneous or shared environments, at the cost of additional computation and memory overhead.
Experimental results on both homogeneous and heterogeneous systems show that, in the worst case, the parallel approaches require up to $1.25\times$ the time of the previous best expansion technique, while enabling shrink operations that are at least $20\times$ faster. In summary, this work provides a practical solution for enabling fast shrink operations without compromising expansion performance.
Future work includes study how to further reduce the synchronization (Subsection 4.3) and connection (Subsection 4.4) overheads, with the aim of outperforming even the basic Merge in expansions. Another promising direction is the design of data redistribution strategies that minimize transfers, ensuring that processes after resizing retain as much of their data as possible.

### v2

@@ 5 Experimental results
The experiments have been designed to address the following questions:
  • Do the proposed parallel strategies reduce the resize time when expanding in comparison with the Merge method?
  • How do these strategies behave in terms of resize time during shrink operations when compared to a TS method?
The first question evaluates whether the proposed strategies provide an improvement over the Merge method. Even if no improvement is observed, it is still necessary to assess whether the additional overhead is justified, given that the primary goal of the parallel strategies is to enable the use of the extremely fast TS method for shrinkage. The second question addresses two objectives. Firstly, it quantifies the difference in resize time between the parallel strategies and the TS method, highlighting the effectiveness of the latter. Secondly, it enables direct comparison with results previously reported in the literature [17].
@@ 5.1 Hardware and Software
The experiments were conducted on two clusters. The first one, Marenostrum 5 (MN5)  [^2] https://www.bsc.es/supportkc/docs/MareNostrum5/overview, used $32$ nodes from the general queue. Each node has two 56-core Intel Xeon 8480 CPU s ($3584$ cores in total). This system was employed to analyze the behavior of both spawning strategies described in Section 4.
@@
The second cluster, Nasp [^3] https://www.hpca.uji.es/computing-facilities/, consists of two sets of 8 nodes. The first set has nodes with two 10-core Intel Xeon 4210 CPU s ($160$ cores total), connected through a 100 Gbit/s InfiniBand Endpoint Detection and Response (EDR) and a 10 Gbit/s Ethernet network. The second set uses nodes with 32-core Intel Xeon 6346 CPU s ($256$ cores total) connected via 10 Gbit/s Ethernet. Both sets communicate through a shared 10 Gbit/s Ethernet link between their switches. This heterogeneous cluster was used to evaluate the effectiveness of the heterogeneous-aware strategy in Section 4.2.
All MN5 experiments were run with MPICH 4.2.0 and CH4:OFI (InfiniBand), while NASP experiments were run with MPICH 3.4.3 and CH3:Nemesis (Ethernet).
Proteo was used to evaluate the different approaches. Before reconfiguration, 5 iterations of Monte Carlo Pi computation including one MPI_Allgather were performed to ensure MPI initialization.
For expansions, MaM was configured using Baseline (B) and Merge (M) methods, each combined with the two strategies described in Section 4. All these configurations are compared against Merge without strategies, which was shown to be the best expansion method in a previous study [17]. For shrinking, MaM was configured to execute B combined with both strategies described in Section 4, and these configurations were compared against the Merge method without strategies. When shrinking, the Merge method in MaM corresponds now to the TS method [^4] Previously, it was the ZS method, as it was not possible to fully terminate zombies.; therefore, since no processes are spawned, it cannot use the described parallel strategies. Nevertheless, thanks to the use of parallel strategies in previous resizes, it ensures that TS will return the nodes to the RMS instead of creating zombies. Moreover, since the ZS method provided the best resizing times in a previous work [17], its counterpart TS is used here both to complement the results of that study and to show the benefits of TS. Each configuration was executed $20$ times, and the median result was reported for every triplet of sources, targets, and method.
@@
Both the Proteo version [^5] https://lorca.act.uji.es/gitlab/martini/malleability_benchmark/-/tree/Paper-ParallelSpawn and the experimental results [15] are publicly available.
5.2 Evaluation on Homogeneous Job Allocation
@@
Both spawning strategies described in Section 4 were evaluated on the homogeneous cluster MN5. In total, $5$ configurations were tested for expansion and $3$ for shrinking. Each configuration involved a single reconfiguration from NS to NT processes, using up to $42$ nodes combinations from the set $\{1,2,4,8,16,24,32\}$.
@@
Figure 4 reports the median reconfiguration times depending on the number of initial nodes ($I$) and resulting nodes ($N$). Figure 4(a) (top) shows expansion results, where the Merge method outperforms all others in $17$ out of $21$ cases ($80.9\%$). Parallel Merge (in both versions) follows closely, requiring up to $1.13\times$ more time due to additional synchronizations among groups. This overhead grows when more than $8$ groups are created and the number of groups is not a power of two. In those cases, at least three steps are required for the binary connection, and the presence of unbalanced leaves increases the cost.
@@
Parallel Baseline methods are consistently slower (up to $1.73\times$), except in one case ($1$ node to $16$ nodes). This overhead stems from spawning extra processes, which also leads to oversubscription.
@@
Figure 4(b) (bottom) shows shrinking results. In this case, avoiding process spawning is clearly advantageous, resulting in a speedup of $1387\times$ compared to the other evaluated approaches. These findings are consistent with previous work [17].
@@
Figure 4: Resizing times in an homogeneous allocation for MN5. The number of processes can be computed by multiplying the nodes by $112$ cores. Letter B refers to the Baseline method, M to Merge method and TS to the Merge shrink variant that terminates ranks.
Figure 5 summarizes the best method for each ($I$, NN) pair based on statistical tests. Axes are defined with Initial Nodes on the vertical and Resulting Nodes on the horizontal; thus, the upper triangle corresponds to expansion and the lower triangle to shrinkage. When multiple methods appear in a cell, they are statistically equivalent, ordered by ascending time.
The main difference arises in expansions with $8$ or fewer groups, requiring at most $3$ steps for the binary connection. In such cases, the parallel methods are preferred as they enable shrinking without spawning processes, while maintaining times comparable to Merge, though statistically higher.
Figure 5: Preferred methods for process management phase depending on the number of nodes $I$ and NN, in MN5. The number of processes can be computed by multiplying the nodes by $112$ cores.
5.3 Evaluation of Iterative Diffusive strategy
The improved strategy described in Section 4.2 was evaluated only on the heterogeneous cluster NASP. Three configurations were tested for expansion and two for shrinkage. In this case, the Hypercube strategy is not included in the evaluation because it is unable to correctly spawn the processes. Each configuration involved a single reconfiguration from NS to NT processes, using up to $72$ combinations of nodes from the set $\{1,2,4,6,8,10,12,14,16\}$. In all cases, the number of nodes of each type was balanced (half of one type and half of the other). When only one node was used, the 20-core node was selected.
Figure 6 shows the median reconfiguration times as a function of initial nodes ($I$) and resulting nodes ($N$). Figure 6(a) (top) reports expansion results, where the Merge method consistently outperforms the others. The iterative diffusive variant of Merge achieves comparable results, with at most a $1.25\times$ increase, but its cost increases slightly with the number of generated groups. Baseline methods are always more expensive, ranging from $1.76\times$ to $4.92\times$ slower, due to oversubscription caused by spawning additional processes.
Figure 6(b) (bottom) presents the shrinking results. As observed in the expansion analysis, the B-based strategies are always the least efficient, as in this case M + TS method achieves speedups of at least $20\times$. To reach these benefits a parallel spawn must be performed beforehand, since TS requires that each MCW to be contained within a single node. Although this parallel spawning for expansions introduces an overhead compared to the previously optimal spawning method [17], the gains in subsequent shrinking operations outweigh that cost.
Furthermore, in $28$ out of $32$ cases ($87.5\%$) for expanding and shrinking, Merge delivered statistically lower times than the iterative diffusive strategy.
@@ 5.3 Evaluation of Iterative Diffusive strategy
Figure 6: Resizing times in an heterogeneous allocation for NASP. The number of processes can be computed by multiplying the nodes by $52$ cores, with the exception for the node, where there are 2020 cores. Letter B refers to the Baseline method, M to Merge method and TS to the Merge shrink variant that terminates ranks.
@@ 6 Conclusions and Future Work
This work introduces and evaluates two new strategies for spawning processes in the traditional MPI. Both strategies rely on parallel spawning and have two main goals: (1) reducing the cost of process creation, and (2) enabling efficient shrink operations using the Merge method, the fastest available option. To achieve the second goal, each MCW must be fully contained within a single node. Previous Merge implementations could not achieve this because processes that should have been terminated during shrinkage remained as zombies preventing node reuse.
Between the two strategies, the Hypercube approach is limited to homogeneous allocations, whereas the Iterative Diffusive strategy can also be used for resource allocations that use nodes with an heterogeneous number of cores.
Experimental results on both homogeneous and heterogeneous allocations on different systems show that, in the worst case, the expansion with parallel approaches require up to $1.25\times$ times against the Merge method, but they enable shrink operations with a TS method that is at least $20\times$ faster than the Baseline method. In summary, this work provides a practical solution for enabling fast shrink operations without compromising expansion performance.
Future work includes studying how to further reduce the synchronization and connection (Subsections 4.3 and 4.4) overheads, with the aim of outperforming even the basic Merge expansions. Additionally, the load balancing challenges associated with heterogeneous hardware have not been considered in this work and must be addressed. Another promising direction is the design of data redistribution strategies that minimize transfers, ensuring that processes retain as much data as possible after resizing.
Acknowledgements
We would like thank Sergio Iserte for his valuable contributions to this work, his insightful discussions on possible improvements of the model, as well as his support in gaining access to the MN5 supercomputing facilities, which enabled the large-scale simulations presented in this article.
@@ Acknowledgements
Data availability
The source codes for MaM are publicly available (https://lorca.act.uji.es/gitlab/martini/malleability_benchmark/-/tree/Paper-ParallelSpawn). All data used to create tables and figures in the Results section are available at [15].
@@
CRediT authorship contribution statement
  • Iker Martín-Álvarez: Conceptualization, Methodology, Software, Investigation, Validation, Formal analysis, Data Curation, Visualization, Writing - Original Draft.
  • José I. Aliaga: Methodology, Writing - Review & Editing, Supervision, Resources, Funding acquisition.
  • Maribel Castillo: Methodology, Formal analysis, Writing - Review & Editing, Supervision.
Declaration of generative AI use
During the preparation of this work the authors used ChatGPT exclusively to refine the language and improve textual clarity. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.